# ReAttention: Training-Free Infinite Context with Finite Attention Scope

### v2

@@ Abstract
The long-context capability of the Large Language Models (LLM) has made significant breakthroughs, but the maximum supported context length remains a critical bottleneck limiting their practical applications. The constraint of context length in LLMs arises from the self-attention mechanism, which cannot effectively and efficiently capture the semantic relationships within infinitely long contexts via the limited pre-trained positional information and attention scope. In this work, we propose ReAttention, a training-free approach enabling LLM based on the self-attention mechanism to support an infinite context with a finite attention scope under sufficient memory resources. ReAttention performs the position-agnostic top-$k$ attention before the ordinary position-aware self-attention, freeing LLMs from the length extrapolation issue. We validate the performance of ReAttention on the LongBench, L-Eval, and InfiniteBench and demonstrate that it is on par with traditional methods. Furthermore, we also apply ReAttention on mainstream LLMs, including LLaMA3.1-8B and Mistral-v0.3-7B, enabling them to support context lengths of at least 1M and even expanding the context length of LLaMA3.2-3B-chat by 128$\times$ to 4M without any further training in Needle-In-A-Haystack tests. We also improve the efficiency of ReAttention with Triton and achieve an efficient extrapolation without additional overhead.
@@ 1 Introduction
Large Language Models (LLM) based on Transformer(Vaswani et al., 2017; OpenAI, 2023; Reid et al., 2024; Sun et al., 2024) has made great progress in the applications of Natural Language Processing (NLP). Particularly, in long context modeling, a considerable amount of research has been dedicated to extending the context length of LLMs(Chen et al., 2023; Peng et al., 2023; Liu et al., 2023; Xiong et al., 2023), increasing the maximum supported context length from the initial 2K(Touvron et al., 2023) to 2M(Ding et al., 2024) in open-source LLMs. However, the maximum context length still remains a bottleneck that limits the practical applications of LLMs(Press et al., 2022; Chen et al., 2023). To achieve infinite context length under sufficient memory for Transformer-based LLMs, the following three conditions must be satisfied:
@@
Early works in extrapolation focus on the first point, primarily by interpolating the position embedding(Chen et al., 2023; bloc97, 2023b; a; Xiong et al., 2023) or limiting the relative position within the pre-training context length(Su, 2023; Jin et al., 2024; An et al., 2024). However, later studies have shown that even if positional information is not OOD, the attention entropy tends to increase logarithmically with the length of the self-attention window(Han et al., 2023). Specifically, the self-attention distribution becomes increasingly diffuse as inference length expands, making it difficult to effectively gather information from the context, which leads to unstable model outputs(Peng et al., 2023; Han et al., 2023). In response, methods like LM-Infinite(Han et al., 2023) and StreamingLLM(Xiao et al., 2023) have been proposed, which retain the starting and ending parts of the input, allowing LLMs to maintain stable outputs as input length increases. However, these approaches compromise the global awareness of contextual information, significantly affecting the performance on downstream tasks.
@@
For human beings, although we have long-term memory, we only need a small amount of information to think and act in real-time. Similarly, while an LLM requires the full context to complete the inference process, it only needs limited contextual information at each inference step. This fact has inspired research like InfLLM(Xiao et al., 2024) and LongHeads(Lu et al., 2024) to perceive and extract critical information from the context before performing self-attention. However, LongHeads still face upper limits on positional information due to the use of NTK or PI for prefill(Lu et al., 2024), while InfLLM suffers from bias in representing contextual information, raised from the chunk-wise representation and ReRoPE-style position embedding when extracting key information.
@@
In this work, we introduce ReAttention, a training-free method that extends LLMs with finite context lengths to process infinite contexts. ReAttention regards extracting critical contextual information as an additional attention process preceding traditional self-attention, akin to ”thinking before you act.” At each reasoning step, ReAttention selects the most critical finite segments from the KV cache without position information, concatenates them, applies positional embedding, and performs self-attention. By controlling the length of the selected KV cache segments, ReAttention can achieve infinite context length with a finite attention scope, while ensuring the position information and attention entropy are not out-of-distribution(Han et al., 2023). Furthermore, inspired by the optimization techniques used in self-attention, particularly FlashAttention (Dao et al., 2022; Dao, 2023), we use Triton (Tillet et al., 2019), a GPU programming language, to minimize read and write overheads in top-$k$ attention. With our custom Triton kernel, ReAttention avoids extra computational overhead and reduces memory usage for long contexts compared to traditional methods. Our contributions can be summarized as follows:
@@
  • We outline three fundamental conditions for infinite context extension in Transformer-based LLMs, position embedding not OOD, stable attention entropy, and effective awareness of critical contextual information. We also find that the third condition can be satisfied through the attention score without positional embedding.
@@
  • We validate that ReAttention matches the performance of self-attention in long-context scenarios with no computational overhead and less memory usage. Specifically, ReAttention extends the context length of leading LLMs, such as LLaMA3.1-8B (Dubey et al., 2024), to at least 1M tokens. For smaller models like LLaMA3.2-3B-chat (Meta, 2024b), the context length can be increased by 128$\times$, reaching up to 4M tokens, without additional training.
@@ 2.1 Full-Context Cache Selection
While LLMs require a complete long context to perform the entire inference process, only a limited context segment is needed at each inference step (Xiao et al., 2023; 2024; Lu et al., 2024). Considering the beginning and end of the input context correspond to globally important prompts and local information for inference (Xiao et al., 2023), ReAttention retains both the global and local segments of the KV cache for inclusion in the self-attention process.
@@
Unlike chunk-based selection in previous works (Lu et al., 2024; Xiao et al., 2024), ReAttention extracts critical contextual information with the dot-product between $\bm{q}_{t}$ and $\bm{\mathcal{K}}_{\text{cache}}$, rather than a chunk-wise representation of $\bm{\mathcal{K}}_{\text{cache}}$. This method offers strong adaptability and avoids the semantic fragmentation caused by fixed chunking (Luo et al., 2024). Moreover, since the intermediate result of $\bm{q}_{t}\bm{\mathcal{K}}_{\text{middle}}^{T}$ is too large for long-context scenarios, and would incur huge read and write overhead, we follow the approach of FlashAttention(Dao et al., 2022; Dao, 2023), fusing the whole process into one kernel using Triton(Tillet et al., 2019), as detailed in Section 3.4.
@@ 2.2 Training-Free Integration
ReAttention concatenates the selected KV cache segments between the global and local parts, applies positional embedding sequentially, and preserves the relative order while ignoring the absolute distance between the selected segments, as shown in Figure 1. Self-attention can then be applied to the concatenated KV cache.
@@
Furthermore, the unselected KV cache segments are unimportant in the current inference step, as their self-attention scores are minimal (Zhang et al., 2024d). Therefore, this modification does not harm the self-attention output and can eliminates interference from irrelevant information(Zhang et al., 2024d; Ge et al., 2023). Compared with InfLLM(Xiao et al., 2024), ReAttention’s position embedding and attention transform within the pre-training window do not introduce the untrained attention pattern, thereby preventing bias accumulation in the KV cache during the prefilling stage. Thus, ReAttention can disregard position information, achieve an unlimited attention context with a limited attention scope without any training, and remain compatible with the existing attention acceleration methods(Dao et al., 2022; Dao, 2023).
@@ 3.1 Setup
We conduct experiments on LLaMA3-8B-8K(Meta, 2024a), LLaMA3.1-8B-128K(Dubey et al., 2024), LLaMA3.1-70B-128K(Dubey et al., 2024), LLaMA3.2-3B-128K(Dubey et al., 2024), Mistral-v0.3-7B-32K(mistralai, 2024), InternLM2.5-7B-1M(InternLM, 2024), Qwen2-7B-128K(Yang et al., 2024), Qwen2-72B-128K(Yang et al., 2024), Qwen2-1B-32K(Yang et al., 2024). For all models, we set the length of $\bm{\mathcal{K}}_{\text{global}}$ to 32, the length of $\bm{\mathcal{K}}_{\text{local}}$ to 4096, and selected span size to 32. Moreover, we set $k=4,k^{\prime}=127$ in top-$k$ attention. Importantly, the attention scope in each step remains within the maximum attention window. For example, for the LLaMA3-8B-8K with ReAttention, the maximum attention scope size is 32 + 4096 + 127$\times$32, which exactly matches the maximum supported attention window of 8192. We use OpenCompass(Contributors, 2023b) for validation. All experiments are performed with FP16 precision and accelerated with FlashAttention2(Dao, 2023).
@@ 3.2 Long-Context Benchmark Evaluation
(a) LongBench En. tasks
(b) LongBench Zh. tasks
(c) LongBench Average score
(d) L-Eval Close-ended tasks
(e) L-Eval Open-ended tasks
(f) L-Eval Average score
Figure 2: Results of existing mainstream LLMs, including LLaMA Series(Meta, 2024a; Dubey et al., 2024; Meta, 2024b) and Mistral(mistralai, 2024) on LongBench(Bai et al., 2023) and L-Eval(An et al., 2023). ReAttention achieves a consistent superiority over StreamingLLM(Xiao et al., 2023) and shows comparable performance with LLMs with full attention.
@@
LLaMA3-8B-Instruct-8K - - - - - - - - - -
+ DynamicNTK 50.66 7.44 21.76 - - - - - - -
@@
+ ReAttention 19.21 11.80 16.75 17.47 11.82 15.06 20.09 11.91 14.13 15.36
Table 1: Results of existing mainstream LLMs, namely LLaMA Series(Meta, 2024a; Dubey et al., 2024; Meta, 2024b) on InfiniteBench(Zhang et al., 2024b) in different context lengths. ”-” means LLM could not provide a stable output in a certain context length. ReAttention achieves superiority over InfLLM(Xiao et al., 2024) and full attention.
We first evaluate all 9 LLMs on the commonly used long-context benchmark LongBench(Bai et al., 2023) and L-Eval(An et al., 2023), with a default context length of 32K and a middle truncation. For LLaMA3-8B-8K, which has a context length of less than 32K, we report its performance at its pre-trained context lengths of 8K and the results with Dynamic NTK(bloc97, 2023a), uesd to extrapolate to the 32K context length. For the Dynamic NTK implementation, we use the default settings from the Huggingface Transformers(Wolf et al., 2020), setting the scaling factor to 4. Additionally, we compare the performance of all 9 LLMs with StreamingLLM(Xiao et al., 2023) as an ablation study, using the same global and local segment lengths as the ReAttention settings.
As shown in Figure 2 and detailed in Table 4, ReAttention outperforms StreamingLLM across all 9 models, indicating that the full-context selection acquires useful information for long-context inference. Furthermore, ReAttention performs on par with full attention and even surpasses it in some cases, such as LLaMA3.1-70B-128K(Meta, 2024a) and Qwen2-1B-32K(Yang et al., 2024). This demonstrates that ReAttention can be applied to LLMs of various sizes and achieve performance close to full attention on downstream tasks(Jiang et al., 2024).
To further demonstrate the superiority and extrapolation capability of ReAttention, we validate our method on InfiniteBench(Zhang et al., 2024c), a more challenging benchmark with a longer context length. We choose 3 commonly tested subtasks, En.MC, En.QA and En.Sum, evaluate models with varying context lengths, and compare ReAttention with DynamicNTK(bloc97, 2023a) and InfLLM(Xiao et al., 2024) with the same selection config. The results [^1] Due to the extensive evaluation for all 9 models, and considering that InfLLM has limited support for larger models such as LLaMA3.1-70B-chat and other series like InternLM2.5 and Qwen2, we only report results of the LLaMA series models at the 8B and 3B scales. are shown in Table 1. Remarkably, ReAttention consistently outperforms full-attention and InfLLM at the 128K context length and in the average score. While DynamicNTK performs well at 32k, it encounters a clear upper bound on extrapolation, beyond which the model fails to produce stable outputs. Moreover, while InfLLM can extend context length indefinitely (Xiao et al., 2024), it still lags behind ReAttention in downstream tasks due to inaccurate extraction of critical information and differences in position embedding format compared to the pre-training phase.
@@ 3.3 Needle-In-A-Haystack Evaluation
Building on this, we increase the difficulty of the evaluation by conducting experiments on multi-NIAH and single-NIAH in longer contexts, with results shown in Figure 4. Given the increased difficulty, we employ the corresponding instruct versions of the models. On the mainstream LLaMA3.1-8B-Instruct-128K(Dubey et al., 2024), we achieve a context length of 1M in the multi-NIAH task. For smaller models, such as LLaMA3.2-3B-chat-128K(Meta, 2024b), we extend the context length to at least 2M. Notably, for Qwen2-1B-Instruct-32k(Yang et al., 2024), we extend the context length to 4M on 4 A100 GPUs, achieving a training-free context expansion of 128$\times$. To the best of our knowledge, it is the largest amplification of context length for LLMs achieved without additional training. This demonstrates that ReAttention can extend the context length while disregarding the position information, using a finite attention window to achieve an infinite attention context.

### v3

@@ Abstract
The long-context capability of the Large Language Models (LLM) has made significant breakthroughs, but the maximum supported context length in length extrapolation remains a critical bottleneck limiting their practical applications. The constraint of context length in LLMs arises from the self-attention mechanism, which cannot effectively and efficiently capture the semantic relationships within infinitely long contexts via the limited pre-trained positional information and attention scope. In this work, we propose ReAttention, a training-free approach enabling LLM based on the self-attention mechanism to support an infinite context with a finite attention scope under sufficient memory resources. ReAttention performs the position-agnostic top-$k$ attention before the ordinary position-aware self-attention, freeing LLMs from the length extrapolation issue. We validate the performance of ReAttention on the LongBench, L-Eval, and InfiniteBench and demonstrate that it is on par with traditional methods. Furthermore, we also apply ReAttention on mainstream LLMs, including LLaMA3.1-8B and Mistral-v0.3-7B, enabling them to support context lengths of at least 1M and even expanding the context length of LLaMA3.2-3B-chat by 128$\times$ to 4M without any further training in Needle-In-A-Haystack tests. We also improve the efficiency of ReAttention with Triton and achieve an efficient extrapolation without additional overhead. The code is available at https://github.com/OpenMOSS/ReAttention.
@@ 1 Introduction
Large Language Models (LLM) based on Transformer (Vaswani et al., 2017; OpenAI, 2023; Reid et al., 2024; Sun et al., 2024) has made great progress in the applications of Natural Language Processing (NLP). Particularly, in long context modeling, a considerable amount of research has been dedicated to extending the length capabilities of LLMs (Chen et al., 2023; Peng et al., 2023; Liu et al., 2023; Xiong et al., 2023), increasing the maximum supported context length from the initial 2K (Touvron et al., 2023) to 2M (Ding et al., 2024) in open-source LLMs. However, the maximum context length in length extrapolation remains a bottleneck that limits the practical applications of LLMs (Press et al., 2022; Chen et al., 2023). To achieve infinite context length under sufficient memory for Transformer-based LLMs, the following three conditions must be satisfied:
@@
Early works in extrapolation focus on the first point, primarily by interpolating the position embedding (Chen et al., 2023; bloc97, 2023b; a; Xiong et al., 2023) or limiting the relative position within the pre-training context length (Su, 2023; Jin et al., 2024; An et al., 2024). However, later studies have shown that even if positional information is not OOD, the attention entropy tends to increase logarithmically with the length of the self-attention window (Han et al., 2023). Specifically, the self-attention distribution becomes increasingly diffuse as inference length expands, making it difficult to effectively gather information from the context, which leads to unstable model outputs (Peng et al., 2023; Han et al., 2023). In response, methods like LM-Infinite (Han et al., 2023) and StreamingLLM (Xiao et al., 2023) have been proposed, which retain the starting and ending parts of the input, allowing LLMs to maintain stable outputs as input length increases. However, they compromise the global awareness of contextual information, harming the downstream performance.
@@
For human beings, although we have long-term memory, we only need a small amount of information to think and act in real-time. Similarly, while an LLM requires the full context to complete the inference process, it only needs limited contextual information at each inference step. This fact has inspired research like InfLLM (Xiao et al., 2024a) and LongHeads (Lu et al., 2024) to perceive and extract critical information from the context before performing self-attention. However, LongHeads still face upper limits on positional information due to the use of NTK or PI for prefill (Lu et al., 2024), while InfLLM suffers from bias in representing contextual information, raised from the chunk-wise representation and ReRoPE-style position embedding when extracting key information.
@@
In this work, we introduce ReAttention, a training-free method that extends LLMs with finite context lengths to process infinite contexts. ReAttention regards extracting critical contextual information as an additional attention process preceding traditional self-attention, akin to ”thinking before you act.” At each reasoning step, ReAttention selects the most critical finite segments from the KV cache without position information, concatenates them, applies positional embedding, and performs self-attention. By controlling the length of the selected KV cache segments, ReAttention can achieve infinite context length with a finite attention scope, while ensuring the position information and attention entropy are not out-of-distribution (Han et al., 2023). Furthermore, inspired by the optimization techniques used in self-attention, particularly FlashAttention  (Dao et al., 2022; Dao, 2023), we use Triton  (Tillet et al., 2019), a GPU programming language, to minimize read and write overheads in top-$k$ attention. With our custom Triton kernel, ReAttention avoids the extra computational overhead and reduces memory usage for long contexts. Our contributions are summarized as follows:
@@
  • We outline three requirements for infinite context in Transformer-based LLMs, position embedding not OOD, stable attention entropy, and effective contextual awareness. We also find that the last one can be satisfied via the attention score without positional embedding.
@@
  • We validate that ReAttention matches the performance of traditional self-attention in long contexts with no computational overhead and less memory usage. Specifically, ReAttention extends the context length of leading LLMs, such as LLaMA3.1-8B-128K, to at least 1 million tokens. For smaller models like LLaMA3.2-3B-chat, the context length can be increased by 128$\times$, reaching up to 4 million tokens, without additional training.
@@ 2.1 Full-Context Cache Selection
While LLMs require a complete long context to perform the entire inference process, only a limited context segment is needed at each inference step (Xiao et al., 2023; 2024a; Lu et al., 2024). Considering the beginning and end of the input context correspond to globally important prompts and local information for inference (Xiao et al., 2023), ReAttention retains both the global and local segments of the KV cache for inclusion in the self-attention process.
@@
Unlike chunk-based selection in previous works  (Lu et al., 2024; Xiao et al., 2024a), ReAttention extracts critical contextual information with the dot-product between $\bm{q}_{t}$ and $\bm{\mathcal{K}}_{\text{cache}}$, rather than a chunk-wise representation of $\bm{\mathcal{K}}_{\text{cache}}$. This method offers strong adaptability and avoids the semantic fragmentation caused by fixed chunking  (Luo et al., 2024). Moreover, since the intermediate result of $\bm{q}_{t}\bm{\mathcal{K}}_{\text{middle}}^{T}$ is too large for long-context scenarios, and would incur huge read and write overhead, we follow the approach of FlashAttention (Dao et al., 2022; Dao, 2023), fusing the whole process into one kernel using Triton (Tillet et al., 2019), as detailed in Section 3.4.
@@ 2.2 Training-Free Integration
ReAttention concatenates the selected KV cache segments between the global and local parts, applies positional embedding sequentially, and preserves the relative order while ignoring the absolute distance between the selected segments, as shown in Figure 1. Self-attention can then be applied to the concatenated KV cache. The pseudocode of the whole process is detailed in Appendix A.
@@
Furthermore, the unselected KV cache segments are unimportant in the current inference step, as their self-attention scores are minimal  (Zhang et al., 2024d). Therefore, this modification does not harm the self-attention output and can eliminates interference from irrelevant information (Zhang et al., 2024d; Ge et al., 2023). Compared with InfLLM (Xiao et al., 2024a), ReAttention’s position embedding and attention transform within the pre-training window do not introduce the untrained attention pattern, thereby preventing bias accumulation in the KV cache during the prefilling stage. Thus, ReAttention can disregard position information, achieve an unlimited attention context with a limited attention scope without any training, and remain compatible with the existing attention acceleration methods (Dao et al., 2022; Dao, 2023).
@@ 3.1 Setup
We conduct experiments on LLaMA3-8B-8K (Meta, 2024a), LLaMA3.1-8B-128K (Dubey et al., 2024), LLaMA3.1-70B-128K (Dubey et al., 2024), LLaMA3.2-3B-128K (Dubey et al., 2024), Mistral-v0.3-7B-32K (mistralai, 2024), InternLM2.5-7B-1M (InternLM, 2024), Qwen2-7B-128K (Yang et al., 2024a), Qwen2-72B-128K (Yang et al., 2024a), Qwen2-1B-32K (Yang et al., 2024a). For all models, we set the length of $\bm{\mathcal{K}}_{\text{global}}$ to 32, the length of $\bm{\mathcal{K}}_{\text{local}}$ to 4096, and selected span size to 32. Moreover, we set $k=4,k^{\prime}=127$ in top-$k$ attention. Importantly, the attention scope in each step remains within the maximum attention window. For example, for the LLaMA3-8B-8K with ReAttention, the maximum attention scope size is 32 + 4096 + 127$\times$32, which exactly matches the maximum supported attention window of 8192. We use OpenCompass (Contributors, 2023b) for validation. All experiments are performed with FP16 precision and accelerated with FlashAttention2 (Dao, 2023).
@@ 3.2 Long-Context Benchmark Evaluation
S-Doc M-Doc Sum ICL Syn. Code Avg.
LLaMA3-8B-8K 18.08 13.17 21.42 63.95 10.45 69.89 35.48
+ StreamingLLM 15.14 8.90 18.36 58.19 7.28 70.27 32.49
+ InfLLM 12.43 8.93 13.20 66.08 7.55 70.03 32.51
+ ReAttention (Ours) 19.08 11.33 20.69 63.40 8.95 70.38 35.03
LLaMA3.1-8B-128K 23.61 15.29 16.15 64.61 31.81 70.63 38.79
+ StreamingLLM 21.67 11.46 14.01 61.05 33.43 70.50 37.06
+ InfLLM 17.11 9.70 15.12 68.46 34.30 68.37 37.14
+ ReAttention (Ours) 23.97 15.22 15.66 64.54 31.20 70.48 38.63
LLaMA3.2-3B-128K 34.45 26.24 24.46 59.29 40.11 55.73 40.76
+ StreamingLLM 28.05 19.41 22.67 55.23 22.63 54.16 35.08
+ InfLLM 17.47 6.77 17.44 49.11 23.63 48.09 28.21
+ ReAttention (Ours) 32.53 24.39 24.13 58.13 37.39 55.17 39.43
Table 1: Results of LLaMA Series(Meta, 2024a; Dubey et al., 2024; Meta, 2024b) on LongBench(Bai et al., 2023). ReAttention achieves a consistent superiority over StreamingLLM(Xiao et al., 2023) and InfLLM(Xiao et al., 2024a) and shows comparable performance with LLMs with full attention.
@@
LLaMA3-8B-Instruct-8K 50.66 7.44 21.76 - - - - - - -
+ StreamingLLM 27.95 4.40 18.41 28.38 4.30 17.61 34.93 4.36 17.64 17.55
@@
+ ReAttention (Ours) 19.21 11.80 16.75 17.47 11.82 15.06 20.09 11.91 14.13 15.36
Table 2: Results of LLaMA Series on InfiniteBench (Zhang et al., 2024b) in different context lengths. ”-” means LLM could not provide a stable output in a certain context length. ReAttention achieves superiority over StreamingLLM (Xiao et al., 2023), InfLLM (Xiao et al., 2024a) and full attention.
We first evaluate all 9 LLMs on the commonly used long-context benchmark LongBench (Bai et al., 2023) and L-Eval (An et al., 2023), with a default context length of 32K and a middle truncation. For LLaMA3-8B-8K, which has a context length of less than 32K, we report its performance with Dynamic NTK (bloc97, 2023a). For the Dynamic NTK implementation, we use the default settings from the Huggingface Transformers (Wolf et al., 2020), setting the scaling factor to 4. Additionally, we compare the performance of all 9 LLMs with StreamingLLM (Xiao et al., 2023) as an ablation study, using the same global and local segment lengths as the ReAttention settings.
As shown in Table 1 and supplemented in Table 5 in Appendix B, ReAttention outperforms StreamingLLM across all 9 models, indicating that the full-context selection acquires useful information for long-context inference. Furthermore, ReAttention performs on par with full attention and even surpasses it in some cases, such as LLaMA3.1-70B-128K (Meta, 2024a) and Qwen2-1B-32K (Yang et al., 2024a). This demonstrates that ReAttention can be applied to LLMs of various sizes and achieve performance close to full attention on downstream tasks (Jiang et al., 2024).
To further demonstrate the superiority and extrapolation capability of ReAttention, we validate our method on InfiniteBench (Zhang et al., 2024c), a more challenging benchmark with a longer context length. We choose 3 commonly tested subtasks, En.MC, En.QA and En.Sum, evaluate models with varying context lengths, and compare ReAttention with DynamicNTK (bloc97, 2023a) and InfLLM (Xiao et al., 2024a) with the same selection config. The results [^1] Due to the extensive evaluation for all 9 models, and considering that InfLLM has limited support for larger models such as LLaMA3.1-70B-chat and other series like InternLM2.5 and Qwen2, we only report results of the LLaMA series models at the 8B and 3B scales. are shown in Table 2. Remarkably, ReAttention consistently outperforms full-attention and InfLLM at the 128K context length and in the average score. While DynamicNTK performs well at 32k, it encounters a clear upper bound on extrapolation, beyond which the model fails to produce stable outputs. Moreover, while InfLLM can extend context length indefinitely  (Xiao et al., 2024a), it still lags behind ReAttention in downstream tasks due to inaccurate extraction of critical information and differences in position embedding format compared to the pre-training phase.
@@ 3.3 Needle-In-A-Haystack Evaluation
Building on this, we increase the difficulty of the evaluation by conducting experiments on multi-NIAH and single-NIAH in longer contexts, with results shown in Figure 3. Given the increased difficulty, we employ the corresponding instruct versions of the models. On the mainstream LLaMA3.1-8B-Instruct-128K (Dubey et al., 2024), we achieve a context length of 1M in the multi-NIAH task. For smaller models, such as LLaMA3.2-3B-chat-128K (Meta, 2024b), we extend the context length to at least 2M. Notably, for Qwen2-1B-Instruct-32k (Yang et al., 2024a), we extend the context length to 4M on 4 A100 GPUs, achieving a training-free context expansion of 128$\times$. To the best of our knowledge, it is the largest amplification of context length for LLMs achieved without additional training. This demonstrates that ReAttention can extend the context length while disregarding the position information, using a finite attention window to achieve an infinite attention context.